# Exploring the Strongly-Interacting Regime of Effective Multi-Body Interactions in a Trapped Ultracold Atom System

### v2

@@ VI CONCLUSIONS AND PROSPECTS
In this work, we reveal the behaviors of strongly-interacting few-body quantum systems comprising of harmonically confined spinless bosons in a wide range of bare two-body interaction with a high-resolution spectroscopy. Strong interactions are induced between the $(n-1)$ |$\ket{g}$-state bosons and a single |$\ket{e}$-state atom, created as a result of the high-resolution laser excitation. It is thus instructive to consider this injected few-body state in the context of the Bose polaron problem, which may provide an indispensable viewpoint for this system. It is straightforward to apply the same methodology to study strongly-interacting regime of spinful atoms of both fermions and bosons, and also with long-range dipolar interactions provided by magnetic atoms and molecules or Rydberg atoms. Systematic investigations of these various systems would give important insights into, for example, how the few-body limit can illuminate the quantum many-body problem, by taking larger values of the number of atoms, which may provide a possible relation to a strongly-interacting bosonic liquid.
@@ Appendix B Sample preparation
In the configuration of our optical lattice system, the trap angular frequencies for atoms in the ground ${{}^{1}}S_{0}$ and excited ${{}^{3}}P_{2}\ (m_{J}=+2)$ states are ${\bm{\omega}}_{g}=2\pi\times(35.9,35.9,35.9)$ kHz and ${\bm{\omega}}_{e}=2\pi\times(44.0,44.0,38.5)$ kHz, respectively, at the final potential depth of $25\,E_{R}$.
@@ Appendix C Details of the data analysis of spectroscopic results
We also obtain $\sigma_{n,\text{StdDev}}$, a weighted standard deviation of the resonance position shifts, given as $\sigma_{n,\text{StdDev}}=\sqrt{\sum_{i}[w_{n,i}\,(x_{\text{shift},n,i}-x_{\text{shift},n})^{2}]/\sum_{i}w_{n,i}}$, to represent the error of the resonance position shifts for every $n$ at a certain magnetic field 111 The use of $\sigma_{n,\text{StdDev}}$ as a representative for the error is justified if $\sigma_{n,\text{StdDev}}$ is not much smaller than $\sigma_{n,\text{ErrProp}}=1/\sqrt{\sum_{i}w_{n,i}}$ calculated from the fitting errors $\sigma_{\text{shift},n,i}$. However, this is not the case for the resonances of $n=4$ and $n=5$ at $B_{0}=232$ mG, $n=3$ at $280$ mG, and $n=5$ at $327$ mG, where we use $\sigma_{n,\text{ErrProp}}$ instead of $\sigma_{n,\text{StdDev}}$ to represent the error. . In addition, the spectrum suffers from long-term laser frequency drifts, which also affects the relative frequency shifts. This effect is compensated for by a linear interpolation with the reference to the resonance position of $n=2$ or $n=1$. Here, we consider the difference between the resonance positions of $n=1$ or $n=2$ and the linear fit to them, as the error $\sigma_{\text{Drift}}$ of the drift compensation. Namely, $\sigma_{\text{Drift}}=\sqrt{\sum_{i}[w^{\prime}_{i}\,(f_{\text{diff},i}-f_{\text{avg}})^{2}]/\sum_{i}w^{\prime}_{i}}$, where $f_{\text{diff},i}$ is the frequency difference between the resonance positions of $n=1$ or $n=2$ and the linear fit to them in each scan indexed as $i$, $f_{\text{avg}}$ is the average of $f_{\text{diff},i}$ weighted by the fitting errors $\sigma^{\prime}_{\text{shift},n,i}$ of the resonance positions of $n=1$ or $n=2$, and $w^{\prime}_{i}=1/\sigma^{\prime 2}_{\text{shift},n,i}$. Note that $\sigma_{\text{Drift}}$ is typically about 1.0–1.5 kHz. In the end, the value calculated with the following formula is used as the total error $\sigma_{\text{total}}$:
@@ Appendix D Scattering length dependence of two-body interaction energy
where $m=2.89\times 10^{-25}$ kg is the mass of 174Yb atoms, $\bm{\omega}_{g(e)}$ are the trap angular frequencies for the ground (excited) state, $a_{eg}$ is the scattering length between them, and $\delta_{\mathrm{reg}}({\bm{r}})\equiv\delta({\bm{r}})\frac{\partial}{\partial|{\bm{r}}|}|{\bm{r}}|$ with $\hbar=1$. The corresponding Lippmann-Schwinger equation below the threshold at $E_{eg}<E_{g}+E_{e}=\sum_{i=x,y,z}\frac{\omega_{gi}+\omega_{ei}}{2}$ reads
@@
$G({\bm{r}}_{g},{\bm{r}}_{e};{\bm{r}}_{g}^{\prime},{\bm{r}}_{e}^{\prime})$ $\equiv\langle{\bm{r}}_{g},{\bm{r}}_{e}|\frac{1}{E_{eg}-\hat{H}_{0}}|{\bm{r}}_{g}^{\prime},{\bm{r}}_{e}^{\prime}\rangle$ (32)
@@
$\chi({\bm{r}})\equiv\lim_{{\bm{r}}_{g},{\bm{r}}_{e}\to{\bm{r}}}\frac{\partial}{\partial|{\bm{r}}_{g}-{\bm{r}}_{e}|}[|{\bm{r}}_{g}-{\bm{r}}_{e}|\psi({\bm{r}}_{g},{\bm{r}}_{e})]$ (33)
@@
$\equiv\lim_{{\bm{r}}_{g},{\bm{r}}_{e}\to{\bm{r}}}\left[G({\bm{r}}_{g},{\bm{r}}_{e};{\bm{r}}^{\prime},{\bm{r}}^{\prime})+\frac{m}{4\pi|{\bm{r}}_{g}-{\bm{r}}_{e}|}\delta({\bm{r}}-{\bm{r}}^{\prime})\right]$ (36)
@@
$\frac{1}{a_{eg}}\sqrt{\frac{\hbar}{m\omega_{eg}}}=\frac{\sqrt{2}\,\Gamma\!\left(-\frac{U_{eg}}{2\hbar\omega_{eg}}\right)}{\Gamma\!\left(-\frac{1}{2}-\frac{U_{eg}}{2\hbar\omega_{eg}}\right)}.$ (44)
The two-body interaction energy $U_{eg}$ assuming an average trap frequency of
@@
is shown in Fig. 6 as a function of $a_{eg}$. The resulting $U_{eg}$ under such equal frequency approximation is found to be in close agreement with the numerical results, which fully take into account the unequal and anisotropic harmonic potentials for two atoms as in the experimental situation of this study. Because the relative difference is less than 1 %, the equal frequency approximation is employed to plot the theoretical curve in Fig. 3(b) of the main text as well as to compute effective multi-body interactions in Sec. III.
Figure 7: Magnetic field variation of the scattering length $a_{eg}$ in the Feshbach resonance. The solid curve denotes a fit with a fitting function $a_{eg}(B)=a_{bg}-\Delta^{\prime}/(B-B_{0})$, with $a_{bg}=0.13(0.16)\,l_{eg}$, $\Delta^{\prime}=-206(45)$ mG, and $B_{0}=398(13)$ mG. The vertical error bars at each magnetic field strength are the same as the horizontal error bars in Fig. 3(b) of the main text.
@@ Appendix F Revisiting the results in previous studies
In this previous study of imaging Mott insulators in a three-dimensional optical lattice with microwave spectroscopy, the measured dependence of the resonance position on the atomic occupation number $n$ is shown in Fig. 2B (circle symbols) of that paper up to $n=5$. The resonance positions corresponding to each $n$ are read from the figure (no readings are available for error bars) and checked to see if the relationships of Eqs. (18) and (20), valid in the case of no effective four- and five-body interactions, are satisfied. As a result, the left- and right-hand sides of Eq. (18) are $-88$ Hz and $-87$ Hz, respectively, and the left- and right-hand sides of Eq. (20) are $-111$ Hz and $-112$ Hz, respectively. This means that both $n=4$ and $n=5$ shifts in this previous study can be explained by considering only up to effective three-body interactions.
@@
In this previous study of performing high-resolution clock spectroscopy for fermionic 87Sr atoms in a three-dimensional optical lattice, the measured dependence of the resonance position up to the atomic occupation number $n=5$ is shown in Fig. 3 (red circles) of that paper. Here, the authors find that the measured values deviate from the calculated values when only two-body interactions are considered, while they are consistent with the perturbative calculation results that consider up to effective three-body interactions. They also mention in the supplementary information that the contribution of effective four-body interactions in their system is so small that they can be ignored, and in Fig. S1 the measured resonance positions up to $n=5$ in the wide range of mean trap depths show ‘good agreement’ with the calculated values considering up to the effective three-body interactions.

### v4

@@ VI CONCLUSIONS AND PROSPECTS
In this work, we reveal the behaviors of strongly-interacting few-body quantum systems comprising of harmonically confined spinless bosons in a wide range of bare two-body interaction with a high-resolution spectroscopy. Strong interactions are induced between the $(n-1)$ |$\ket{g}$-state bosons and a single |$\ket{e}$-state atom, created as a result of the high-resolution laser excitation. It is straightforward to apply the same methodology to study strongly-interacting regime of spinful atoms of both fermions and bosons, and also with long-range dipolar interactions provided by magnetic atoms and molecules or Rydberg atoms. Systematic investigations of these various systems would give important insights into, for example, how the few-body limit can illuminate the quantum many-body problem, by taking larger values of the number of atoms, which may provide a possible relation to a strongly-interacting bosonic liquid.
@@ Appendix B Sample preparation
In the configuration of our optical lattice system, the trap angular frequencies for atoms in the ground ${{}^{1}}S_{0}$ and excited ${{}^{3}}P_{2}\ (m_{J}=+2)$ states are ${\bm{\omega}}_{g}=2\pi\times(35.9,35.9,35.9)$ kHz and ${\bm{\omega}}_{e}=2\pi\times(44.0,44.0,38.5)$ kHz, respectively, at the final potential depth of $25\,E_{R}$. Note that the inhomogeneity of trap frequencies is estimated to be about 0.2 %.
@@ Appendix C Details of the data analysis of spectroscopic results
We also obtain $\sigma_{n,\text{StdDev}}$, a weighted standard deviation of the resonance position shifts, given as $\sigma_{n,\text{StdDev}}=\sqrt{\sum_{i}[w_{n,i}\,(x_{\text{shift},n,i}-x_{\text{shift},n})^{2}]/\sum_{i}w_{n,i}}$, to represent the error of the resonance position shifts for every $n$ at a certain magnetic field 111 The use of $\sigma_{n,\text{StdDev}}$ as a representative for the error is justified if $\sigma_{n,\text{StdDev}}$ is not much smaller than $\sigma_{n,\text{ErrProp}}=1/\sqrt{\sum_{i}w_{n,i}}$ calculated from the fitting errors $\sigma_{\text{shift},n,i}$. However, this is not the case for the resonances of $n=4$ and $n=5$ at $B_{0}=232$ mG, $n=3$ at $280$ mG, and $n=5$ at $327$ mG, where we use $\sigma_{n,\text{ErrProp}}$ instead of $\sigma_{n,\text{StdDev}}$ to represent the error. . In addition, the spectrum suffers from long-term laser frequency drifts, which also affect the relative frequency shifts. This effect is compensated for by a linear interpolation with the reference to the resonance position of $n=2$ or $n=1$. Here, we consider the difference between the resonance positions of $n=1$ or $n=2$ and the linear fit to them, as the error $\sigma_{\text{Drift}}$ of the drift compensation. Namely, $\sigma_{\text{Drift}}=\sqrt{\sum_{i}[w^{\prime}_{i}\,(f_{\text{diff},i}-f_{\text{avg}})^{2}]/\sum_{i}w^{\prime}_{i}}$, where $f_{\text{diff},i}$ is the frequency difference between the resonance positions of $n=1$ or $n=2$ and the linear fit to them in each scan indexed as $i$, $f_{\text{avg}}$ is the average of $f_{\text{diff},i}$ weighted by the fitting errors $\sigma^{\prime}_{\text{shift},n,i}$ of the resonance positions of $n=1$ or $n=2$, and $w^{\prime}_{i}=1/\sigma^{\prime 2}_{\text{shift},n,i}$. Note that $\sigma_{\text{Drift}}$ is typically about 1.0–1.5 kHz. In the end, the value calculated with the following formula is used as the total error $\sigma_{\text{total}}$:
@@ Appendix D Scattering length dependence of two-body interaction energy
where $m=2.89\times 10^{-25}$ kg is the mass of 174Yb atoms, $\bm{\omega}_{g(e)}$ are the trap angular frequencies for the ground (excited) state, $a_{eg}$ is the scattering length between them, and $\delta_{\mathrm{reg}}({\bm{r}})\coloneqq\delta({\bm{r}})\frac{\partial}{\partial|{\bm{r}}|}|{\bm{r}}|$ with $\hbar=1$ throughout this paragraph. The corresponding Lippmann-Schwinger equation below the threshold at $E_{eg}<E_{g}+E_{e}=\sum_{i=x,y,z}\frac{\omega_{gi}+\omega_{ei}}{2}$ reads
@@
$G({\bm{r}}_{g},{\bm{r}}_{e};{\bm{r}}_{g}^{\prime},{\bm{r}}_{e}^{\prime})$ $\coloneqq\langle{\bm{r}}_{g},{\bm{r}}_{e}|\frac{1}{E_{eg}-\hat{H}_{0}}|{\bm{r}}_{g}^{\prime},{\bm{r}}_{e}^{\prime}\rangle$ (32)
@@
$\chi({\bm{r}})\coloneqq\lim_{{\bm{r}}_{g},{\bm{r}}_{e}\to{\bm{r}}}\frac{\partial}{\partial|{\bm{r}}_{g}-{\bm{r}}_{e}|}[|{\bm{r}}_{g}-{\bm{r}}_{e}|\psi({\bm{r}}_{g},{\bm{r}}_{e})]$ (33)
@@
$\coloneqq\lim_{{\bm{r}}_{g},{\bm{r}}_{e}\to{\bm{r}}}\left[G({\bm{r}}_{g},{\bm{r}}_{e};{\bm{r}}^{\prime},{\bm{r}}^{\prime})+\frac{m}{4\pi|{\bm{r}}_{g}-{\bm{r}}_{e}|}\delta({\bm{r}}-{\bm{r}}^{\prime})\right]$ (36)
@@
$\frac{1}{a_{eg}}\sqrt{\frac{\hbar}{m\omega_{eg}}}=\frac{\sqrt{2}\,\Gamma\!\left(-\frac{U_{eg}}{2\hbar\omega_{eg}}\right)}{\Gamma\!\left(-\frac{1}{2}-\frac{U_{eg}}{2\hbar\omega_{eg}}\right)},$ (44)
whereas anisotropic cases have been studied in Refs. [35, 36, 37, 38]. The two-body interaction energy $U_{eg}$ assuming an average trap frequency of
@@
is shown in Fig. 6 as a function of $a_{eg}$. The resulting $U_{eg}$ under such equal frequency approximation is found to be in close agreement with the numerical results, which fully take into account the unequal and anisotropic harmonic potentials for two atoms as in the experimental situation of this study. Because the relative difference is less than 1 %, the equal frequency approximation is employed to plot the theoretical curve in Fig. 3(b) of the main text as well as to compute effective multi-body interactions in Sec. III. We also note that the contribution of the effective range compared with the dominant term in the $s$-wave scattering amplitude is estimated to be smaller than 1/6 for the scattering length we are working on [55, 56], so that the effective range is neglected in our analysis.
@@ Appendix E Derivation of linear relationships between frequency shifts
Figure 7: Magnetic field variation of the scattering length $a_{eg}$ in the Feshbach resonance. The solid curve denotes a fit on all plots with a fitting function $a_{eg}(B)=a_{bg}-\Delta^{\prime}/(B-B_{0})$, with $a_{bg}=0.13(0.16)\,l_{eg}$, $\Delta^{\prime}=-206(45)$ mG, and $B_{0}=398(13)$ mG. The vertical error bars at each magnetic field strength are the same as the horizontal error bars in Fig. 3(b) of the main text.
@@ Appendix F Revisiting the results in previous studies
In this previous study of imaging Mott insulators in a three-dimensional optical lattice with microwave spectroscopy, the measured dependence of the resonance position on the atomic occupation number $n$ is shown in Fig. 2B (circle symbols) of Ref. [30] up to $n=5$. The resonance positions corresponding to each $n$ are read from the figure (no readings are available for error bars) and checked to see if the relationships of Eqs. (18) and (20), valid in the case of no effective four- and five-body interactions, are satisfied. As a result, the left- and right-hand sides of Eq. (18) are $-88$ Hz and $-87$ Hz, respectively, and the left- and right-hand sides of Eq. (20) are $-111$ Hz and $-112$ Hz, respectively. This means that both $n=4$ and $n=5$ shifts in this previous study can be explained by considering only up to effective three-body interactions.
@@
In this previous study of performing high-resolution clock spectroscopy for fermionic 87Sr atoms in a three-dimensional optical lattice, the measured dependence of the resonance position up to the atomic occupation number $n=5$ is shown in Fig. 3 (red circles) of Ref. [33]. Here, the authors find that the measured values deviate from the calculated values when only two-body interactions are considered, while they are consistent with the perturbative calculation results that consider up to effective three-body interactions. They also mention in the supplementary information that the contribution of effective four-body interactions in their system is so small that they can be ignored, and in Fig. S1 the measured resonance positions up to $n=5$ in the wide range of mean trap depths show ‘good agreement’ with the calculated values considering up to the effective three-body interactions.